OpenStudy (anonymous):

let l = longg s= short 4l=5s-3 15=l+ s Use substitution and solve for the variables

OpenStudy (anonymous):

4L=5s-3 15=L+s solve for any variable in the 2nd equation since its the easiest one. for this I'll pick s s=15-L and now substitute s=15-L into 4L=5s-3 where s is. so it'll look like this 4L=5(15-L)-3 and now just solve for L and you'll get 8. now just plug in L=8 in the equation 15=L+s and solve for s. You get 7. so s=7 and L=8